OpenStudy (elise_a18):

Factor completely 2x3 + 4x2 + 6x + 12. 2(x3 + 2x2 + 3x + 6) (2x2 + 6)(x + 2) 2[(x2 + 2)(x + 3)] 2[(x2 + 3)(x + 2)]

OpenStudy (anonymous):

2x^3 + 4x^2 + 6x + 12 = 2 (x^3 + 2x^2 + 3x + 6) = 2 ( ( x^3 + 3x ) + (2x^2 + 6 ) ) = 2 ( x ( x^2 + 3 ) + 2 (x^2 + 3 ) ) = 2 ( x + 2 ) ( x^2 + 3 ) @Elise_a18

OpenStudy (elise_a18):

A?